1/1

William Wordsworth,

Evening on Calais Beach

It is a beauteous evening, calm and free, The holy time is quiet as a Nun Breathless with adoration; the broad sun Is sinking down in its tranquillity;

The gentleness of heaven broods o’er the sea: Listen! the mighty Being is awake, And doth with his eternal motion make A sound like thunder—everlastingly.

Dear Child! dear Girl! that walkest with me here, If thou appear untouch’d by solemn thought, Thy nature is not therefore less divine:

Thou liest in Abraham’s bosom all the year; And worshipp’st at the Temple’s inner shrine, God being with thee when we know it not.

Verkinto de tiu ĉi Angla poemo estas William Wordsworth (∗1770 – †1850).

Arg-312-630 (2005-05-19 10:27:47)

http://www.poezio.net/version?poem-id=312&version-id=630